Aaron plays the saxophone but his neighbor Bruno hates music. Aaron and Bruno have the following utility functions: UA = 100 + 10z - 0.1z2 and UB = 100 - 10z; where z is the number of hours spent by Aaron playing the saxophone (and UA includes the cost of the saxophone). Determine:

a) [3 marks] The number of hours of saxophone played by Aaron when the external effect on Bruno is ignored.

b) [3 marks] The socially optimal level of hours of saxophone that should be played by Aaron.

c) [5 marks] The optimal Pigouvian tax needed to reach the social optimum.

d) [7 marks] The outcome with Coasian bargaining when the property right is assigned to the saxophone player. [Hint: Calculate the utility gain of Aaron and the utility loss of Bruno. What is the minimum amount of money Aaron needs as compensation? How much Bruno is willing to pay to reduce the number of hours of saxophone?].

e) [7 marks] The outcome with Coasian bargaining when the property right is assigned to the neighbor.
Answer #1
  • Answers for : (a.) 50  (b.) 0 (c.) 10
  • answers for rest and explanation is given below :

UA= 100 + 102 0.1 22 UR = 100 - 102 the utility maximization Aaron Sawes paoblem : V 12) = 100 + 10 2 - 0.12? max {2>,o} TheThis function is decreasing in z , Assume Aaron has to pay a tax © for each hour. The utility maximization problem is now max@ When the property right is assigned to saxophone player , the chooses to play . hours. The neighbour (Bruno)s utility lo sassigned to ban When the property sight is to neighbour, the chooses Saxophone. The Aarons utility loss is! VA (0) - VA (50)